# SVarM: Linear Support Varifold Machines for Classification and Regression on Geometric Data

### v1

@@ 3 Model Architecture
The core of our regression model architecture is to approximate a map $h:\mathbb{R}^{n}\times S^{n-1}\to\mathbb{R}$ (for $n=2$ or $3$) which can itself approximate any affine map on $\mathcal{V}$ as shown in Theorem LABEL:thm:affine_approx. We perform this approximation using a multilevel perceptron (MLP) $h_{\theta}:\mathbb{R}^{2n}\to\mathbb{R}$ where $\theta$ denotes the trainable parameters of the MLP. It is then straightforward, based on the many existing universal approximation and expressivity theorems [25, 17, 24, 32], to rephrase all the results of Section 2.3 for functions $h$ that are expressible by a sufficiently wide and/or deep MLP.
@@
where $\beta$ is a trainable bias term. For the linear regression problem in $\mathcal{V}$, we then train this model in a supervised manner by which, given a set of varifolds $\{\mu_{1},...,\mu_{R}\}$ (computed from the corresponding input shapes) and the associated labels $\{y_{1},...,y_{R}\}$, we minimize the mean squared error loss
@@ 4.1 Regression for Single Axis Rotation Alignment of Human Bodies
To address this, we frame rotational alignment as a regression problem, where the model learns to predict the angle by which a given mesh has been rotated. The goal is to realign each mesh to a canonical upright position. We generate our dataset by applying known rotations to a subset of the FAUST 3D human body scans along a single axis, storing the corresponding angles as labels resulting in  8000 labeled meshes. Our model is trained on  90% of this data and evaluated on the remaining  10%. In Figure 4, we present a selection of dataset samples and demonstrate the model’s ability to accurately predict rotation angles and realign the meshes.
@@ 4.3 Classification of MNIST Digits
We next evaluate SVarM on the task of classifying handwritten digits from the well-known MNIST dataset. To adapt this image dataset to our framework, following a similar setup to [29], we convert each digit into a 3D surface representation by treating the grayscale pixel values as a height map. We then apply a remeshing operation to generate closed surface meshes corresponding to each digit. A schematic of the conversion process, along with representative examples of the resulting surface meshes, is shown in Figure 6. This pre-processing is not computationally burdensome and requires $<$20 minutes for the entire MNIST dataset.
@@
4.4 Classification of Faces from the COMA Shape Graph Dataset
To further evaluate the effectiveness for classification, we train SVarM on a dataset of shape graphs derived from the CoMA dataset [41], which contains 3D surface scans of human faces. This dataset comprises approximately 21,000 scans of 12 distinct individuals. From each scan, we extract a shape graph by identifying a consistent set of facial landmarks and connecting them via geodesic paths along the surface of the mesh. A schematic illustration of this shape graph construction process is provided in Figure 7. We should note that shape graphs, just as curves, are naturally embedded in the space of varifolds by simple addition in $\mathcal{V}$ of the varifolds associated to each separate branch.
@@ 4.5 Robustness Studies
5 Conclusion and Limitations
@@
In this work, we propose a learning framework on datasets of curves, surfaces and shape graphs that leverages the (nonlinear) embedding of shapes into varifold space. We focus on this particular measure space as it provides a parameterization-invariant representation of embedded shapes, making the model particularly well-suited for geometric datasets. Analogous to support vector machines—but in the infinite-dimensional setting of varifolds—the SVarM model then learns affine forms $\mu\mapsto\langle\mu,h\rangle+\beta$ to solve regression and classification problems involving 3D geometric data, by approximating a test function $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1},\mathbb{R})$ and bias $\beta$. We derive some theoretical guarantees on SVarM’s capacity to approximate arbitrary affine forms, identify conditions under which sets of varifolds or embedded shapes can be linearly separated, and establish stability results for the model. In addition, we present a series of numerical experiments demonstrating the model’s effectiveness in both regression and classification tasks across various geometric datasets. These experiments highlight the model’s computational efficiency and its robustness to imaging noise, including missing faces and sampling changes. The results also showcase the fact that SVar can achieve comparable (and in some cases superior) accuracy to geometric deep learning models but with an overall much smaller number of trainable parameters, making it in line with recent efforts towards developing lightweight deep learning architectures [31]. Lastly, although the present work was focused on objects such as curves and surfaces, the versatility of the varifold representation makes it possible, in principle, to extend SVarM to other types of geometric data structures, with two cases of interest being high-angular resolution diffusion images [47] and spatial transcriptomics data [45] as these naturally fit the varifold setting.
Limitations: In spite of the above advantages, the SVarM model has certain limitations that we aim to address in future work. In particular, while we demonstrate that any affine function on the space of varifolds can be approximated, the model is not capable of representing non-linear functions on this space. As a result, although we showed how affine functions on $\mathcal{V}$ may be enough to separate disjoint sets of shapes under relatively mild assumptions, there still exist configurations that cannot be classified using this model, c.f. counterexample of Figure 2. In future work, we aim to investigate the learning of non-linear functions on the space of varifolds by combining multiple affine functions with non-linear activation functions, similar to a two-layer multilayer perceptron architecture. This could prove useful in tackling other types of problems than shape classification and regression. Incidentally, we hope to establish a universal approximation theorem for continuous functions on the space of varifolds. Another interesting aspect that has not yet been explored in this paper is the use, during training, of regularization penalties on the weights of the neural network defining $h$. Such penalties could enforce certain constraints on $h$ so as to potentially increase the qualitative interpretability of the learned test function.

### v2

@@ 3.2 Neural network model
As an alternative to the above kernel strategy, we will instead consider the more direct approach of modeling the test function via a neural network, which is now ubiquitous in the modern machine learning literature. We shall thus approximate $h$ using a multilevel perceptron (MLP), namely consider a set of trainable functions $h_{\theta}:\mathbb{R}^{2n}\to\mathbb{R}$ where $\theta$ denotes the trainable parameters of the MLP. It is then straightforward, based on the many existing universal approximation and expressivity theorems [28, 19, 27, 37], to rephrase all the results of Section 2.3 for functions $h$ that are expressible by a sufficiently wide and/or deep MLP.
@@
where $\beta$ is a trainable bias term. For the linear regression problem in $\mathcal{V}$, we train this model in a supervised manner. Similar to the previous kernel approach, we minimize the mean squared error loss to the known labels:
@@
Remark 3 (Invariance to Orientation).
It is important to note that the model presented so far relies on the orientation of the tangent/normals to the shapes, meaning that one assumes a consistent orientation of the different shapes in the dataset. For certain applications that involve non-oriented data, it is necessary to make the SVarM approach invariant to the orientation of $v_{i}$. This can easily be achieved by taking $\widehat{h_{\theta}}=h_{\theta}(x,v)+h_{\theta}(x,-v)$ and computing $\langle\mu,\widehat{h_{\theta}}\rangle+\beta$. In particular, in the applications to shape graphs of Section 4.4, we rely on this modification to remain robust to reorientation of the edges of the shape graphs.
@@ 4.1 Regression for Single Axis Rotation Alignment of Human Bodies
To address this, we frame rotational alignment as a regression problem, where the model learns to predict the angle by which a given mesh has been rotated. The goal is to realign each mesh to a canonical upright position. We generate our dataset by applying known rotations to a subset of the FAUST 3D human body scans along a single axis, storing the corresponding angles as labels resulting in  8000 labeled meshes. Our model is trained on  90% of this data and evaluated on the remaining  10%. In Figure 4, we present a selection of dataset samples and demonstrate the model’s ability to accurately predict rotation angles and realign the meshes. Moreover, to highlight the interpretability of our model, we display the evaluation of the learned test function $h$ on samples from the dataset where the color of each face of the mesh corresponds to the evaluation of the test function at $(c_{f},n_{f})$. We note that the meshes that are misaligned by a larger angle have a higher density of faces with larger values of $h$.
@@ 4.2 Regression for Full Rotation Alignment of Human Data
To further highlight our method’s effectiveness at solving rotation alignment problems we compare with results attained by directly optimizing over the rotation group to minimize the varifold distance (some explanation probably required here). This approach requires no training time but requires 15-30s per mesh to solve the alignment problem. Directly solving the optimization for each element of the testing set produces a mean error of $8.0558^{\circ}$ with an $R^{2}$ value of $0.799$. By comparison, our trained model requires less than 15s to align the entire testing set and produces a mean error of $3.147^{\circ}$ with an $R^{2}$ value of $0.986$.
@@ 4.3 Classification of MNIST Digits
We next evaluate SVarM on the task of classifying handwritten digits from the well-known MNIST dataset. To adapt this image dataset to our framework, following a similar setup to [34], we convert each digit into a 3D surface representation by treating the grayscale pixel values as a height map. We then apply a remeshing operation to generate closed surface meshes corresponding to each digit. A schematic of the conversion process, along with representative examples of the resulting surface meshes, is shown in Figure 6. This pre-processing is not computationally burdensome and requires <<20 minutes for the entire MNIST dataset.
@@
4.4 Classification of Shape Graphs
To further evaluate the effectiveness of the SVarM model for classification tasks, we train the model on two separate datasets of shape graphs. First, we train SVarM on a dataset of shape graphs derived from the CoMA dataset [46], which contains 3D surface scans of human faces. This dataset is comprised of approximately 21,000 scans of 12 distinct individuals. From each scan, we extract a shape graph by identifying a consistent set of facial landmarks and connecting them via geodesic paths along the surface of the mesh. A schematic illustration of this shape graph construction process is provided in Figure 7. We should note that shape graphs, just as curves, are naturally embedded in the space of varifolds by simple addition in $\mathcal{V}$ of the varifolds associated to each separate branch.
@@ 4.4 Classification of Shape Graphs
Figure 8: Neuron Shape Graphs. Examples of Neuron shape graphs from five different regions of of the mouse nervous system. The second row displays the learned test function for each class in the same coordinates orientation as in the first row. Color represents the average of the function at a given spatial location over the spherical component, while black curves on each pixel illustrate the direction profile of the function at that point. Lastly, in the third row, we show the overall spherical profile of each test function averaged over all spatial points within the image range.
The second dataset consists of 3D shape graphs representing neurons from five distinct regions of the mouse nervous system. This dataset includes 443 neuron graphs, categorized into the Basal Ganglia, Cerebellum, Hippocampus, Main Olfactory Bulb, and Retina. Representative examples from each class are shown in Figure 8. We train the model on  80% neuron graphs and evaluate its performance on the remaining  20%. The trained model correctly classifies 97.73% of the shape graphs in the testing set and Figure 8 shows the spatial and directional profiles of the estimated test functions for each class. This experiment highlights the ability of the SVarM model to be trained on smaller datasets without overfitting. In particular, this model is over-parameterized with 1525 trainable parameters and achieves similar accuracy on the training data as on the unseen data.
@@ 4.5 Robustness Studies
5 Towards nonlinear regressors and classifiers in $\mathcal{V}$
In spite of the mathematical and numerical results presented in this paper, one of the obvious limitations of the SVarM model is that it is restricted to regression and classification functionals that are affine on the space of varifolds. Although we showed that these may be enough to separate disjoint sets of shapes under relatively mild assumptions, the counterexample of Figure 2 still illustrates that some simple configurations are not separable within the SVarM framework. A natural way of generalizing the approach would be thus to consider nonlinear functionals on $\mathcal{V}$, by analogy with the machine learning models based on neural network representations that are commonly used for data living on Euclidean spaces.
In the context of the infinite-dimensional space of varifolds $\mathcal{V}$, we may build nonlinear maps $\psi:\mathcal{V}\rightarrow\mathbb{R}$ via the composition of multiple affine forms on $\mathcal{V}$ with some nonlinear function on $\mathbb{R}$. More specifically, a simple single hidden layer architecture would consist in modeling $\psi$ as follows:
$\psi(\mu)=\sum_{k=1}^{r}c_{k}\sigma(\langle\mu,h_{k}\rangle+\beta_{k})$ (14)
[…]
Theorem 6.
Let $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ be a continuous non-polynomial function, $C$ a weak-* compact subset of $\mathcal{V}$ and $\psi:C\rightarrow\mathbb{R}$ a weak-* continuous function on $C$. Then for any $\varepsilon>0$, there exist $r\in\mathbb{N}$, $(h_{k})_{k=1,\ldots,r}$ in $C_{0}(\mathbb{R}^{n}\times S^{n-1})$, $(\beta_{k})\in\mathbb{R}^{r}$ and $(c_{k})\in\mathbb{R}^{r}$ such that for all $\mu\in C$:
$\left|\psi(\mu)-\sum_{k=1}^{r}c_{k}\sigma(\langle\mu,h_{k}\rangle+\beta_{k})\right|\leq\varepsilon$
The statement follows directly from the more general result of Theorem 2.1 in [30] since the space $\mathcal{V}$ equipped with the weak-* convergence topology is a locally convex space which continuous linear forms are exactly of the form $\mu\mapsto\langle\mu,h\rangle$ (c.f. Proposition 1). Combined with our earlier remarks in Section 3.2, this implies that it is possible to approximate any weak-* continuous (nonlinear) function on $\mathcal{V}$ by considering sufficiently many test functions each represented with sufficiently wide (or deep) MLPs.
Using such more general architectures should theoretically enable to address, in particular, the nonseparability issues of SVarM. For instance, it is easy to derive a function $\psi$ of the form (14) with only one test function $h$ that can separate the pairs of shapes in the example of Figure 2. Namely, if $h$ is defined so that $\langle\mu,h\rangle$ corresponds to the number of balls on the upper half of the object and $\sigma(x)=2^{-|x-1|^{2}}$, then $\psi=1$ for the green surfaces while $\psi=1/2$ for the red ones. We leave it to future work, however, to implement and experiment these types of architectures on shape regression and classification problems.
6 Conclusion and future perspectives
@@
In this work, we propose a learning framework on datasets of curves, surfaces and shape graphs that leverages the (nonlinear) embedding of shapes into varifold space. We focus on this particular measure space as it provides a parameterization-invariant representation of embedded shapes, making the model particularly well-suited for geometric datasets. Analogous to support vector machines—but in the infinite-dimensional setting of varifolds—the SVarM model then learns affine forms $\mu\mapsto\langle\mu,h\rangle+\beta$ to solve regression and classification problems involving 3D geometric data, by approximating a test function $h\in C_{0}(\mathbb{R}^{n}\times S^{n-1},\mathbb{R})$ and bias $\beta$. We derive some theoretical guarantees on SVarM’s capacity to approximate arbitrary affine forms, identify conditions under which sets of varifolds or embedded shapes can be linearly separated, and establish stability results for the model. In addition, we present a series of numerical experiments demonstrating the model’s effectiveness in both regression and classification tasks across various geometric datasets. These experiments highlight the model’s computational efficiency and its robustness to imaging noise, including missing faces and sampling changes. The results also showcase the fact that SVar can achieve comparable (and in some cases superior) accuracy to geometric deep learning models but with an overall much smaller number of trainable parameters, making it in line with recent efforts towards developing lightweight deep learning architectures [36].
While we focused on learning affine functions, we also outlined a more general approach to build and learn nonlinear operators over the space of varifolds, which we plan to implement and experiment for various types of supervised or unsupervised problems involving shape datasets in the future. Furthermore, although the present work was primarily concerned with objects such as curves and surfaces, the versatility of the varifold representation makes it possible, in principle, to extend SVarM to other types of geometric data structures, with two cases of interest being high-angular resolution diffusion images [52] and spatial transcriptomics data [50] as these naturally fit the varifold setting. Lastly, an interesting aspect that has not yet been explored in this paper is the use, during training, of regularization penalties on the weights of the neural network defining $h$. Such penalties could enforce certain constraints on $h$ so as to potentially increase the qualitative interpretability of the learned test function.
Acknowledgments
This work was supported by the National Science Foundation through the grants DMS-2438562 and DMS-2426550. The authors would also like to thank Murad Hossein for his assistance with the mouse neuron traces dataset used in the numerical experiments section.